# Nonstabilizerness Estimation using Graph Neural Networks

### v1

@@ Abstract
This article proposes a Graph Neural Network (GNN) approach to estimate nonstabilizerness in quantum circuits, measured by the stabilizer Rényi entropy (SRE). Nonstabilizerness is a fundamental resource for quantum advantage, and efficient SRE estimations are highly beneficial in practical applications. We address the nonstabilizerness estimation problem through three supervised learning formulations starting from easier classification tasks to the more challenging regression task. Experimental results show that the proposed GNN manages to capture meaningful features from the graph-based circuit representation, resulting in robust generalization performances achieved across diverse scenarios. In classification tasks, the GNN is trained on product states and generalizes on circuits evolved under Clifford operations, entangled states, and circuits with higher number of qubits. In the regression task, the GNN significantly improves the SRE estimation on out-of-distribution circuits with higher number of qubits and gate counts compared to previous work, for both random quantum circuits and structured circuits derived from the transverse-field Ising model. Moreover, the graph representation of quantum circuits naturally integrates hardware-specific information. Simulations on noisy quantum hardware highlight the potential of the proposed GNN to predict the SRE measured on quantum devices.
@@ 1 Introduction
Quantifying the computational resources required to achieve a quantum advantage is a main problem in quantum information theory. Nonstabilizerness, also known as magic [bravyi2005universal, campbell2017roads], emerges as a key resource for quantum advantage as it measures the deviation of a quantum state from the polynomial-time simulability in the framework of the Gottesman-Knill Theorem [gottesman1998heisenberg, aaronson2004improved]. This theorem states that any quantum computation that consists exclusively of Clifford operations can be perfectly simulated in polynomial time by a classical machine. However, preparing a generic quantum state in the full Hilbert space requires resources beyond the Clifford group, since a universal gate set cannot be constructed using Clifford gates alone. The amount of these non-Clifford resources is referred to as nonstabilizerness.
@@
To rigorously capture the effect of non-Clifford operations in quantum computations, several measures of nonstabilizerness have been proposed. Examples include negative quasi-probability [veitch2012negative], robustness of magic [howard2017application, heinrich2019robustness], stabilizer nullity [beverland2020lower], Bell magic [haug2023scalable], and stabilizer Rényi entropies [leone2022stabilizer, leone2024stabilizer]. Among these measures, we focus on the Stabilizer Rényi Entropy (SRE) [leone2022stabilizer] for two main reasons. First, SRE offers favorable computational properties [leone2022stabilizer, leone2024stabilizer, ahmadi2024quantifying], especially when compared to alternative measures that require costly optimization procedures [veitch2012negative, howard2017application, heinrich2019robustness]. Second, SRE is particularly suitable for experiments on real quantum devices [oliviero2022measuring, niroula2024phase]. However, in the general case, the computational time required to calculate the SRE of a quantum state scales exponentially in the number of qubits [leone2022stabilizer]. Tensor Network (TN) methods are currently the most widely adopted approach to calculate SRE for quantum states [haug2023quantifying, tarabunga2024nonstabilizerness, tarabunga2023many, lami2023nonstabilizerness, lami2024unveiling]. Although TN methods offer a theoretical framework with polynomial scaling, they are practically limited to low bond dimension, which restricts the class of quantum states that can be efficiently represented. Recently, Neural Quantum States [sinibaldi2025non] have been proposed as a promising alternative to overcome these limitations by flexibly approximating quantum states with high entanglement. Motivated by previous research in machine learning applications to quantum technology [carleo2019machine, dunjko2018machine, krenn2023artificial], learning-based approaches have recently been proposed for trading off online computational cost with approximate nonstabilizerness estimation. On the one hand, in the restricted case of the stabilizer state classification, where the task is to distinguish between stabilizer and magic states [mello2025retrieving], Convolutional Neural Networks achieve predictions that are robust under the Clifford group. On the other hand, classical machine learning approaches trained on tabular data achieve accurate estimation of SRE, but with limited generalization performances [lipardi2025study].
@@
In this article, we formulate the nonstabilizerness estimation as a graph-level prediction task and propose a Graph Neural Network (GNN) [scarselli2008graph] approach to bridge the gap between the bond dimension-limited TN methods and the limited performances of machine learning models. We test the performance of the proposed GNN approach through three different formulations of nonstabilizerness estimation with increasing difficulty by exploiting the capabilities of GNN to capture meaningful features in the graph-based circuit representation.
@@
    We propose a GNN approach for the nonstabilizerness estimation problem, that exhibits higher generalization capabilities to larger quantum circuits, in terms of number of qubits and gate counts, and to unseen entanglement, compared to previous works [mello2025retrieving, lipardi2025study]. The experimental results highlight the positive impact of representing quantum circuits as graphs, which enables the model to effectively capture their intrinsic structural properties.
@@
    We have generated and released a comprehensive dataset of quantum circuits tailored to address the nonstabilizerness estimation problem in the supervised learning setting [lipardi2025dataset]. The dataset contains a total of $1686000$ quantum circuits designed to span a broad range of number of qubits, entanglement and SRE values. The dataset is structured to address the problem in three different formulations of increasing difficulty, including both classification and regression tasks. We expect this dataset to foster progress in the development and benchmarking of machine learning methods for nonstabilizerness estimation.
@@ 2.1 Stabilizer States
Let us consider an $n$-qubit quantum state and let $U(n)$ be the group of unitary operators on the Hilbert space $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$. We define the Pauli group $\mathcal{\widetilde{P}}_{n}$ acting on $\mathcal{H}$ as the set of all possible Pauli strings $\mathcal{P}_{n}$ taken with multiplicative factors $\pm 1$ and $\pm i$, and the Clifford group $\mathcal{C}(n)\subset U(n)$ as the normalizer of the Pauli group [nielsen2010quantum]:
@@
As proved by the Gottesman-Knill Theorem [gottesman1998heisenberg], stabilizer states can be perfectly simulated by a classical computer in polynomial time. As a consequence, quantum computations restricted to stabilizer states and Clifford gates do not provide any computational advantage over the classical counterpart. Because the set of stabilizer states is closed under Clifford operations, and a universal quantum gate set cannot be constructed from Clifford gates alone, the preparation of a generic state in the Hilbert space requires resources beyond the Clifford group. The amount of such non-Clifford resources is referred to as the nonstabilizerness of the state.
Since it is possible to prepare fully-entangled quantum states that can be efficiently simulated on a classical computer  [nielsen2010quantum, aaronson2004improved], entanglement is not the only resource to consider in the quest for quantum advantage. Nonstabilizerness, also known as magic, plays a fundamental role in assessing quantum advantage [gottesman1998heisenberg, bravyi2005universal, campbell2017roads], and, as such, it is important to define a measure for it. A measure of nonstabilizerness should be a nonstabilizer monotone [leone2024stabilizer], invariant under Clifford operations, and ideally efficiently computable for practical applications.
@@
The first measures of nonstabilizerness that have been proposed involve optimization procedures [veitch2012negative, howard2017application, heinrich2019robustness], which come with high computational cost. Then, more tractable measures have been introduced, such as stabilizer nullity [beverland2020lower], Bell magic [haug2023scalable], and stabilizer Rényi entropies [leone2022stabilizer]. In this article, we focus on the stabilizer Rényi entropy (SRE) [leone2022stabilizer] due to its favorable computational properties [haug2023quantifying, tarabunga2023many, lami2023nonstabilizerness, lami2024unveiling, ahmadi2024quantifying], especially when compared to alternative measures that require costly optimization procedures [veitch2012negative, howard2017application, heinrich2019robustness], and its practical suitability for experimental measurements on quantum hardware [oliviero2022measuring, niroula2024phase].
2.2 Stabilizer Rényi Entropy
For a pure $n$-qubit quantum state $\rho$, the Stabilizer Rényi Entropy (SRE) [leone2022stabilizer] of order $\alpha$ (also called Rényi index) is defined as:
where $\Xi_{P}(\rho)=\frac{1}{2^{n}}Tr(\rho P)^{2}$ and $\mathcal{P}_{n}$ denotes the set of $n$-qubit Pauli strings. For Rényi index $\alpha\geq 2$, the $M_{\alpha}$ fulfills all the requirements of a proper measure of nonstabilizerness [leone2024stabilizer, haug2023scalable]. In this article, we fix the index to the commonly used $\alpha=2$ [leone2024stabilizer], and we will refer to SRE simply as $M_{2}$.
Methods based on Matrix Product States [haug2023quantifying, lami2023nonstabilizerness, lami2024unveiling, tarabunga2024nonstabilizerness], and more in general Tensor Networks, offer a framework to compute $M_{2}$ efficiently in the number of qubits but in running time that depends exponentially on the bond dimensions. Then, these methods represent a powerful theoretical framework, but with practical limitations to low bond dimensions [tarabunga2023many]. To overcome those limitations, recent advances have been made towards efficient $M_{2}$ computations using Neural Quantum States [sinibaldi2025non].
In the general case, the computational cost to compute $M_{2}$ of a quantum state using the Equation 4 grows exponentially with respect to the number of qubits $n$ [leone2022stabilizer], as it requires estimating $4^{n}$ expectation values, corresponding to all possible combinations of Pauli strings.
@@ 2.3 Problem Formulation
In this section, we define three formulations of the nonstabilizerness estimation problem on quantum circuits, with the goal of providing tasks of increasing difficulty that can serve as benchmarks for our study and future research. The first is a classification task aimed at distinguishing quantum circuits that prepare stabilizer states from those that prepare magic states. This formulation relies uniquely on the definition of stabilizer states provided in Section 3 and it is therefore independent from all measures of nonstabilizerness, including $M_{2}$. We refer to this task as the stabilizer state classification. The second is a classification task based on SRE [leone2022stabilizer], where the goal is to distinguish between quantum circuits with low and high $M_{2}$ values. The difficulty of this task is parameterized by the $M_{2}$ threshold, which determines the boundary between low-$M_{2}$ and high-$M_{2}$ quantum circuits. We refer to this task as the SRE-based classification. Finally, the third formulation is a regression task, where the goal is to estimate the degree of nonstabilizerness of a quantum circuit by predicting its positive real-valued $M_{2}$. We refer to this task as the SRE estimation.
@@
We emphasize that our formulation targets nonstabilizerness estimation at the circuit level, rather than at the level of quantum states. As a consequence, it does not require any form of quantum state tomography and relies uniquely on the circuit-based quantum circuit representation. On the one hand, the circuit-based formulation offers significant advantages for applications that are inherently defined within the circuit-model paradigm of quantum computing. An example are Variational Quantum Algorithms [cerezo2021variational], whose link to $M_{2}$ has been recently discussed in [spriggs2025quantum]. On the other hand, the circuit-based formulation is not directly applicable to scenarios where the $M_{2}$ must be evaluated on unknown quantum states, for which a corresponding circuit is generally unavailable.
@@
The stabilizer state classification, particularly relevant in the context discussed in [mello2025retrieving], becomes efficiently solvable in its matrix or circuit formulation [de2025fast]. Differently from the stabilizer state classification, the SRE-based classification and the SRE estimation both involve evaluating the contribution of each gate and its parameters to the overall circuit $M_{2}$, depending on their positions, which is generally hard as discussed in Section 2.2. The SRE-based classification is not only computationally difficult but also physically relevant. In experiments on real quantum hardware, quantum noise can affect the measured $M_{2}$ values depending on both the circuit structure and the hardware specifics. By setting an appropriate threshold close to zero, the SRE-based classification can be used to identify stabilizer states in the presence of quantum noise. Lastly, the SRE estimation is the more general and challenging formulation of the nonstabilizerness estimation problem.
@@ 3 Related Work
In the last decade, different review articles have emphasized the growing influence of machine learning in the physical sciences, see for example [carleo2019machine] and the reference therein, particularly within the rapidly evolving field of quantum technologies [dunjko2018machine, krenn2023artificial].
@@
In quantum computing, particular efforts have been made in tackling problems such as fidelity estimation [zhang2021direct, wang2022torchquantum, saravanan2022data, vadali2024quantum], quantum compiling [acampora2021deep, zlokapa2020deep, zhang2020topological, fan2022optimizing, ruiz2025quantum], ground-state energy estimation in many-body quantum systems [lewis2024improved], quantum error mitigation [liao2024machine], quantum circuit expressibility estimation [aktar2024graph]. Following this research line, supervised learning approaches have been recently proposed to estimate the nonstabilizerness of quantum states [mello2025retrieving, lipardi2025study]. The first is a Convolutional Neural Network (CNN) approach proposed to distinguish between stabilizer and magic states [mello2025retrieving]. In this work, the authors construct a balanced dataset of stabilizer and magic states restricted to product states of fixed size, with $18$ qubits. Each state is represented by a $\mathcal{N}\times n$ binary matrix, where each row corresponds to a single-shot measurement outcome in the computational basis. This matrix is then converted into an image by assigning white pixels to zeros and black pixels to ones, enabling the use of CNN architectures. The CNN model achieves promising results, including predictions that are robust under the Clifford group, which means good generalization capabilities on entangled states. However, the CNN approach is limited to the stabilizer state classification problem and to fixed size of quantum states, because the number of qubits affects the size of the images. The second approach extended the nonstabilizerness estimation problem to a regression task for the supervised machine learning setting [lipardi2025study]. For this purpose, they use the $M_{2}$ as a measure of nonstabilizerness for quantum circuits. In the general case, computing $M_{2}$ of quantum states scales exponentially with the qubit number. The authors tackle the SRE estimation problem with different machine learning models, including Random Forest Regressor and Support Vector Regressor, providing a study on various tabular classical representations of quantum circuits, including circuit-level feature and expectation values of $1$ and $2$-local observables calculated using classical shadows [huang2020predicting]. The Support Vector Regressor (SVR) trained on feature vectors encoding the gate counts achieve the best performance. Altough this approach showed promising experimental results to trade off online computational cost for approximate $M_{2}$ with machine learning, the models exhibit limited generalization capabilities to unseen larger circuits, in terms of both number of qubits and total gate count.
@@
In this article, we advance beyond both the CNN and SVR approaches by proposing a Graph Neural Network (GNN) approach for nonstabilizerness estimation. Our method naturally accommodates both classification and regression formulations of the problem, while offering improved generalization capabilities compared to previous works [mello2025retrieving, lipardi2025study]. It is important to note that, in general, more complex models such as GNNs do not necessarily guarantee better performance. For example, in a recent study on supervised machine learning approaches to quantum error mitigation, Random Forests were shown to outperform other models, including GNNs [liao2024machine]. The choice of data representations and model architectures is often task-dependent, and the design of the appropriate combination of features and models facilitate the learning of relevant properties and ensure robust generalization performance.

### v2

@@ Abstract
This article proposes a Graph Neural Network (GNN) approach to estimate nonstabilizerness in quantum circuits, measured by the stabilizer Rényi entropy (SRE). Nonstabilizerness is a fundamental resource for quantum advantage, and efficient SRE estimations are highly beneficial in practical applications. We address the nonstabilizerness estimation problem through three supervised learning formulations starting from easier classification tasks to the more challenging regression task. Experimental results show that the proposed GNN manages to capture meaningful features from the graph-based circuit representation, resulting in robust generalization performances achieved across diverse scenarios. In classification tasks, the GNN is trained on product states and generalizes on circuits evolved under Clifford operations, entangled states, and circuits with higher number of qubits. In the regression task, the GNN significantly improves the SRE estimation on out-of-distribution circuits with higher number of qubits and gate counts compared to previous work, for both unstructured random quantum circuits and structured circuits derived from the transverse-field Ising model. Moreover, the graph representation of quantum circuits naturally integrates hardware-specific information. Simulations on noisy quantum hardware highlight the potential of the proposed GNN to predict the SRE measured on quantum devices.
@@ 1 Introduction
Quantifying the computational resources required to achieve a quantum advantage is a main problem in quantum information theory. Nonstabilizerness, also known as magic [6, 7], emerges as a key resource for quantum advantage as it measures the deviation of a quantum state from the polynomial-time simulability in the framework of the Gottesman-Knill Theorem [16, 1]. This theorem states that any quantum computation that consists exclusively of Clifford operations can be perfectly simulated in polynomial time by a classical machine. However, preparing a generic quantum state in the full Hilbert space requires resources beyond the Clifford group, since a universal gate set cannot be constructed using Clifford gates alone. The amount of these non-Clifford resources is referred to as nonstabilizerness.
@@
To rigorously capture the effect of non-Clifford operations in quantum computations, several measures of nonstabilizerness have been proposed. Examples include negative quasi-probability [52], robustness of magic [23, 22], stabilizer nullity [5], Bell magic [17], and stabilizer Rényi entropies [29, 28]. Among these measures, we focus on the Stabilizer Rényi Entropy (SRE) [29] for two main reasons. First, SRE offers favorable computational properties [29, 28, 3], especially when compared to alternative measures that require costly optimization procedures [52, 23, 22]. Second, SRE is particularly suitable for experiments on real quantum devices [40, 39]. However, in the general case, the computational time required to compute the SRE of a quantum state scales exponentially in the number of qubits [29]. Tensor Network (TN) methods are currently the most widely adopted approach to compute SRE for quantum states [18, 49, 50, 26, 27]. Although TN methods offer a theoretical framework with polynomial scaling, they are practically limited to low bond dimension, which restricts the class of quantum states that can be efficiently represented. Recently, Neural Quantum States [46] have been proposed as a promising alternative to overcome these limitations by flexibly approximating quantum states with high entanglement. Motivated by previous research in machine learning applications to quantum technology [8, 14, 25], learning-based approaches have recently been proposed for trading off online computational cost with approximate nonstabilizerness estimation. On one hand, in the restricted case of the stabilizer state classification, where the task is to distinguish between stabilizer and magic states [36], Convolutional Neural Networks achieve predictions that are robust under the Clifford group. On the other hand, classical machine learning approaches trained on tabular data achieve accurate estimation of SRE, but with limited generalization performances [32].
@@
In this article, we formulate the nonstabilizerness estimation as a graph-level prediction task and propose a Graph Neural Network (GNN) [45] approach to bridge the gap between machine learning models that achieve robust performance on the restricted classification task and models that are capable of estimating SRE on the general regression task but with limited generalization performances.
@@
    We propose a GNN approach for the nonstabilizerness estimation problem, that exhibits higher generalization capabilities to larger quantum circuits, in terms of number of qubits and gate counts, and to unseen entanglement, compared to previous works [36, 32]. The experimental results highlight the positive impact of representing quantum circuits as graphs, which enables the model to effectively capture their intrinsic structural properties.
@@
    We have generated and released a comprehensive dataset of quantum circuits tailored to address the nonstabilizerness estimation problem in the supervised learning setting [34]. The dataset contains a total of $1686000$ quantum circuits designed to span a broad range of number of qubits, entanglement and SRE values. The dataset is structured to address the problem in three different formulations of increasing difficulty, including both classification and regression tasks. We expect this dataset to foster progress in the development and benchmarking of machine learning methods for nonstabilizerness estimation.
@@
We note that the proposed approach trades off the quality of SRE predictions for a significantly reduced prediction time [32]. In fact, after a one-time training procedure, the GNN can approximate SRE for a quantum circuit almost in real time. As such, it should be regarded as a complementary approach to TN methods rather than a competing one. This approach is relevant in applications where an exact computation of $M_{2}$ can be traded off for a faster approximation. Examples include exploring the relationship between favorable solutions in variational quantum algorithms and their corresponding SRE values [48], as well as enabling quantum architecture search methods to guide the circuit design towards quantum circuits with high SRE values.
@@ 2.1 Stabilizer States
Let us consider an $n$-qubit quantum state and let $U(n)$ be the group of unitary operators on the Hilbert space $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$. We define the Pauli group $\mathcal{\widetilde{P}}_{n}$ acting on $\mathcal{H}$ as the set of all possible Pauli strings $\mathcal{P}_{n}$ taken with multiplicative factors $\pm 1$ and $\pm i$, and the Clifford group $\mathcal{C}(n)\subset U(n)$ as the normalizer of the Pauli group [38]:
@@
As proved by the Gottesman-Knill Theorem [16], stabilizer states can be perfectly simulated by a classical computer in polynomial time. As a consequence, quantum computations restricted to stabilizer states and Clifford gates do not provide any computational advantage over the classical counterpart. Because the set of stabilizer states is closed under Clifford operations, and a universal quantum gate set cannot be constructed from Clifford gates alone, the preparation of a generic state in the Hilbert space requires resources beyond the Clifford group. The amount of such non-Clifford resources is referred to as the nonstabilizerness of the state.
Since it is possible to prepare fully-entangled quantum states that can be efficiently simulated on a classical computer [38, 1], entanglement is not the only resource to consider in the quest for quantum advantage. Nonstabilizerness, also known as magic, plays a fundamental role in assessing quantum advantage [16, 6, 7], and, as such, it is important to define a measure for it. A proper measure of nonstabilizerness has to be a nonstabilizer (magic) monotone, as rigorously studied in magic-state resource theory [10, 21, 35].
The first measures of nonstabilizerness that have been proposed involve optimization procedures [52, 23, 22], which come with high computational cost. Then, more tractable measures have been introduced, such as stabilizer nullity [5], Bell magic [17], and stabilizer Rényi entropies [29]. In this article, we focus on the stabilizer Rényi entropy (SRE) [29] due to its favorable computational properties [18, 50, 26, 27, 3], especially when compared to alternative measures that require costly optimization procedures [52, 23, 22], and its practical suitability for experimental measurements on quantum hardware [40, 39].
2.2 Stabilizer Rényi Entropy
For a pure $n$-qubit quantum state $\rho$, the Stabilizer Rényi Entropy (SRE) [29] of order $\alpha$ (also called Rényi index) is defined as:
@@ 2.2 Stabilizer Rényi Entropy
Methods based on Matrix Product States [18, 26, 27, 49], and more in general Tensor Networks, offer a framework to compute $M_{2}$ efficiently in the number of qubits but in running time that scales as a polynomial of degree four in the bond dimension [18]. These methods represent a powerful theoretical framework, but with practical limitations to low bond dimensions [50]. To overcome those limitations, recent advances have been made towards efficient $M_{2}$ computations using Neural Quantum States [46]. However, in the general case, the computational cost to compute $M_{2}$ of a quantum state using the Equation 4 grows exponentially with respect to the number of qubits $n$ [29], as it requires estimating $4^{n}$ expectation values, corresponding to all possible combinations of Pauli strings.
The SRE generalization to mixed states $\tilde{M_{2}}$ is defined by rescaling $M_{2}$ with the logarithm of the purity of the mixed quantum state $\rho$ [29, 28, 40]:
$\tilde{M_{2}}(\rho)=M_{2}(\rho)-\ln Tr(\rho^{2}).$ (5)
We note that $\tilde{M_{2}}$ is zero on the stabilizer mixed state defined as the equally-weighted mixed state over all the $n$-qubits stabilizer states, but it might be greater than zero for some mixtures of stabilizer states. Since $\tilde{M}$ is not a proper nonstabilizer monotone for mixed states, we also consider the witness of magic $\mathcal{W}_{2}$ [20], which satisfy the relation:
$\mathcal{W}_{2}(\rho)=\tilde{M}_{2}(\rho)+2\ln Tr(\rho^{2}).$ (6)
A positive value of $\mathcal{W}_{2}({\rho})$ certifies that the state $\rho$ lies outside the stabilizer polytope, e.g. it is a nonstabilizer state. However, $\mathcal{W}_{2}({\rho})\leq 0$ does not guarantee that $\rho$ is a stabilizer state [20].
We refer the interested reader to the existing literature for further details on stabilizer monotones for mixed states [29, 19, 28, 54, 20, 47].
@@ 2.3 Problem Formulation
In this section, we define three formulations of the nonstabilizerness estimation problem on quantum circuits, with the goal of providing tasks of increasing difficulty that can serve as benchmarks for our study and future research. The first is a classification task aimed at distinguishing quantum circuits that prepare stabilizer states from those that prepare magic states. This formulation relies uniquely on the definition of stabilizer states and it is therefore independent from all measures of nonstabilizerness, including $M_{2}$. We refer to this task as the stabilizer state classification. The second is a classification task based on SRE [29], where the goal is to distinguish between quantum circuits with low and high $M_{2}$ values. The difficulty of this task is parameterized by the $M_{2}$ threshold, which determines the boundary between low-$M_{2}$ and high-$M_{2}$ quantum circuits. We refer to this task as the SRE-based classification. Finally, the third formulation is a regression task, where the goal is to estimate the degree of nonstabilizerness of a quantum circuit by predicting its positive real-valued $M_{2}$. We refer to this task as the SRE estimation.
@@
We emphasize that our formulation targets nonstabilizerness estimation at the circuit level, rather than at the level of quantum states. As a consequence, it does not require any form of quantum state tomography and relies uniquely on the circuit-based quantum circuit representation. On the one hand, the circuit-based formulation offers significant advantages for applications that are inherently defined within the circuit-model paradigm of quantum computing. An example are Variational Quantum Algorithms [9], whose link to $M_{2}$ has been recently discussed in [48]. On the other hand, the circuit-based formulation is not directly applicable to scenarios where the $M_{2}$ must be evaluated on unknown quantum states, for which a corresponding circuit is generally unavailable.
@@
The stabilizer state classification, particularly relevant in the context discussed in [36], becomes efficiently solvable in its matrix or circuit formulation [12]. Differently from the stabilizer state classification, the SRE-based classification and the SRE estimation both involve evaluating the contribution of each gate and its parameters to the overall circuit $M_{2}$, depending on their positions, which is generally hard. The SRE-based classification is not only computationally difficult but also physically relevant. In experiments on real quantum hardware, quantum noise can affect the measured $M_{2}$ values depending on both the circuit structure and the hardware specifics. By setting an appropriate threshold close to zero, the SRE-based classification can be used to identify stabilizer states in the presence of quantum noise. Lastly, the SRE estimation is the more general and challenging formulation of the nonstabilizerness estimation problem.
@@ 3 Related Work
In the last decade, different review articles have emphasized the growing influence of machine learning in the physical sciences, see for example [8] and the reference therein, particularly within the rapidly evolving field of quantum technologies [14, 25].
@@
In quantum computing, particular efforts have been made in tackling problems such as fidelity estimation [56, 53, 44, 51], quantum compiling [2, 58, 57, 15, 43], ground-state energy estimation in many-body quantum systems [30], quantum error mitigation [31], quantum circuit expressibility estimation [4]. Following this research line, supervised learning approaches have been recently proposed to estimate the nonstabilizerness of quantum states [36, 32]. The first is a Convolutional Neural Network (CNN) approach proposed to distinguish between stabilizer and magic states [36]. In this work, the authors construct a balanced dataset of stabilizer and magic states restricted to product states of fixed size, with $18$ qubits. Each state is represented by a $\mathcal{N}\times n$ binary matrix, where each row corresponds to a single-shot measurement outcome in the computational basis. This matrix is then converted into an image by assigning white pixels to zeros and black pixels to ones, enabling the use of CNN architectures. The CNN model achieves promising results, including predictions that are robust under the Clifford group, which means good generalization capabilities on entangled states. However, the CNN approach is limited to the stabilizer state classification problem and to fixed size of quantum states, because the number of qubits affects the size of the images. The second approach extended the nonstabilizerness estimation problem to a regression task for the supervised machine learning setting [32]. For this purpose, they use the $M_{2}$ as a measure of nonstabilizerness for quantum circuits. In the general case, computing $M_{2}$ of quantum states scales exponentially with the qubit number. The authors tackle the SRE estimation problem with different machine learning models, including Random Forest Regressor and Support Vector Regressor, providing a study on various tabular classical representations of quantum circuits, including circuit-level feature and expectation values of $1$ and $2$-local observables calculated using classical shadows [24]. The Support Vector Regressor (SVR) trained on feature vectors encoding the gate counts achieve the best performance. Altough this approach showed promising experimental results to trade off online computational cost for approximate $M_{2}$ with machine learning, the models exhibit limited generalization capabilities to unseen larger circuits, in terms of both number of qubits and total gate count.
@@
In this article, we advance beyond both the CNN and SVR approaches by proposing a Graph Neural Network (GNN) approach for nonstabilizerness estimation. Our method naturally accommodates both classification and regression formulations of the problem, while offering improved generalization capabilities compared to previous works [36, 32]. It is important to note that, in general, more complex models such as GNNs do not necessarily guarantee better performance. For example, in a recent study on supervised machine learning approaches to quantum error mitigation, Random Forests were shown to outperform other models, including GNNs [31]. The choice of data representations and model architectures is often task-dependent, and the design of the appropriate combination of features and models facilitate the learning of relevant properties and ensure robust generalization performance.